# Origen: Zero-Shot 3D Orientation Grounding in Text-to-Image Generation

### v2

@@ 2 Related Work
For controlling image generation, ControlNet [28] and IP-Adapter [47] are commonly used to utilize a pretrained generative model to control for various conditions, though they require training data. For 3D orientation grounding, no public real-world training data is available, and collecting such data would be particularly challenging, especially for multi-object grounding, due to the need for diversity. To address this, recent works have introduced RL-based reward fine-tuning approaches [48, 49, 50, 51, 52]. However, these methods require substantial computational resources and, more importantly, cause significant deviation from the pretrained data distribution when trained with task-specific rewards, leading to degraded image quality and reduced diversity [53]. Hence, instead of fine-tuning, we propose a test-time reward-guided framework that leverages a discriminative foundational model for guidance.
@@ 3.1 Problem Definition and Overview
[…]
@@
In particular, given a reward function $\mathcal{R}$ defined based on GroundingDINO and OrientAnything (which we will discuss in detail in Sec. 3.2), our goal is to find a latent sample $\mathbf{x}$ that maximizes the reward of its corresponding image, expressed as $\mathcal{R}(\mathcal{F}_{\theta}(\mathbf{x},c))$. For simplicity, we define the pullback of the reward function as ^$\hat{\mathcal{R}}=\mathcal{R}\circ\mathcal{F}_{\theta}$ and use this notation throughout.
@@ 3.3 Reward-Guided Langevin Dynamics
Reward-Guided Langevin Dynamics. Let $q=\mathcal{N}(\mathbf{0},\mathbf{I})$ denote the prior distribution, ^$\hat{\mathcal{R}}(\mathbf{x})$ be the pullback of a differentiable reward function, and $\mathbf{w}_{t}$ denote the standard Wiener process. As $t\rightarrow\infty$, the stationary distribution of the following Langevin dynamics
@@ 3.4 Reward-Adaptive Time Rescaling
2:$\text{Initialize }\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\mathcal{R}_{*}=-\inf$
3:for $i=0$ to $M-1$ do
4:    ^$\hat{\mathcal{R}}_{i}=\mathcal{R}(\mathcal{F}_{\theta}(\mathbf{x}_{i},c),\mathcal{W},\Phi)$triangleright Reward Calc.
5:    $\gamma_{i}=\mathcal{G}(\hat{\mathcal{R}}_{i})\gamma$triangleright Timestep rescaling
@@
7:    $\mathbf{x}_{i+1}=\sqrt{1-\gamma_{i}}\mathbf{x}_{i}+\gamma_{i}\eta\nabla\hat{\mathcal{R}}_{i}$$+\frac{1}{2}\gamma_{i}\nabla\log\mathcal{G}(\hat{\mathcal{R}}_{i})$
8:               + $\sqrt{\gamma_{i}}\epsilon_{i}$triangleright Update step
9:    if ^$\hat{\mathcal{R}}_{i}>\mathcal{R}_{*}$ then
10:        =^$\mathcal{R}_{*}=\hat{\mathcal{R}}_{i},\mathbf{x}_{*}=\mathbf{x}_{i}$
@@ 4.2 Evaluation Metrics
We compare Origen’s performance from two perspectives: (1) Orientation Alignment and (2) Text-to-Image Alignment. We measure orientation alignment using two metrics: 1) Acc.@$22.5^{\circ}$, the angular accuracy within a tolerance of ±$\pm 22.5^{\circ}$ and 2) the absolute error on azimuth angles [^1] We perform comparisons only on azimuth angle, as existing methods [18, 17] do not support the control over polar and rotation angles. Note that our results for all azimuth, polar, and rotation angles are provided in Appendix G.1. between the predicted and grounding object orientations, following Wang et al. [20]. For evaluation, we use OrientAnything [20] to predict the 3D orientation from the generated images. Since its predictions may not be perfect, we also conduct a user study in Sec. 4.6 to validate the results. Along with grounding accuracy, we also evaluate text-to-image alignment using CLIP Score [66], VQA-Score [67], and PickScore [68].
@@ 4.7 Computation Time
In our experiments, we set the number of iterations in our method to match the denoising steps in multi-step generative models (e.g., 50), resulting in comparable computation time—approximately 52.7 seconds per image. A detailed comparison is provided in Appendix I. Notably, for cases achieving angular accuracy within a tolerance of ±$\pm 22.5^{\circ}$ (Tab. 3), the average number of iterations to reach this threshold was 12.8, corresponding to an average of 13.5 seconds. Our approach can also serve as a post-refinement method for existing models, potentially further reducing computation time.
@@ Appendix A Analysis on Norm-Based Regularization
where ^$\hat{\mathcal{R}}=\mathcal{R}\circ\mathcal{F}_{\theta}$ is the pullback of the reward function $\mathcal{R}$ and $y(\|\cdot\|)$ represents the probability density of a $\chi^{d}$ distribution.
We can view both $\hat{\mathcal{R}}(\cdot)$ and $y(\|\cdot\|)$ as components of a single reward, allowing us to rewrite the update as
@@ C.1 Non-Convergence of Direct Time-Rescaled SDE
~$\tilde{\mathcal{L}}^{*}q^{*}(\tilde{\mathbf{x}})$ $=-\nabla\cdot\Big{[}\mathcal{G}(\tilde{\mathbf{x}})\Big{(}-\frac{1}{2}\tilde{\mathbf{x}}+\eta\,\nabla\hat{\mathcal{R}}(\tilde{\mathbf{x}})\Big{)}q^{*}(\tilde{\mathbf{x}})\Big{]}+\frac{1}{2}\nabla^{2}\Big{[}\mathcal{G}(\tilde{\mathbf{x}})q^{*}(\tilde{\mathbf{x}})\Big{]}$
$=-\nabla\cdot\Big{[}a(\tilde{\mathbf{x}})\,q^{*}(\tilde{\mathbf{x}})\Big{]}+\frac{1}{2}\nabla\cdot\Big{[}\nabla\mathcal{G}(\tilde{\mathbf{x}})\,q^{*}(\tilde{\mathbf{x}})+\mathcal{G}(\tilde{\mathbf{x}})\,\nabla q^{*}(\tilde{\mathbf{x}})\Big{]}$
$=-\nabla\cdot\Big{[}a(\tilde{\mathbf{x}})\,q^{*}(\tilde{\mathbf{x}})\Big{]}+\frac{1}{2}\nabla\cdot\Big{[}\nabla\mathcal{G}(\tilde{\mathbf{x}})\,q^{*}(\tilde{\mathbf{x}})+2a(\tilde{\mathbf{x}})\,q^{*}(\tilde{\mathbf{x}})\Big{]}$
$=\frac{1}{2}\nabla\cdot\Big{[}\nabla\mathcal{G}(\tilde{\mathbf{x}})\,q^{*}(\tilde{\mathbf{x}})\Big{]}$
@@ C.2 Time-Rescaled SDE with Correction Drift Term
The corresponding Fokker-Planck operator ~$\tilde{\mathcal{L}}^{*}_{\text{corr}}$ for this corrected time-rescaled SDE now annihilates the original stationary distribution $q^{*}(\mathbf{x})$:
~$\tilde{\mathcal{L}}^{*}_{\text{corr}}q^{*}(\tilde{\mathbf{x}})$ $=-\nabla\cdot\Big{[}\mathcal{G}(\tilde{\mathbf{x}})\Big{(}-\frac{1}{2}\tilde{\mathbf{x}}+\eta\,\nabla\hat{\mathcal{R}}(\tilde{\mathbf{x}})\Big{)}q^{*}(\tilde{\mathbf{x}})+\frac{1}{2}\nabla\mathcal{G}(\tilde{\mathbf{x}})\,q^{*}(\tilde{\mathbf{x}})\Big{]}+\frac{1}{2}\nabla^{2}\Big{[}\mathcal{G}(\tilde{\mathbf{x}})q^{*}(\tilde{\mathbf{x}})\Big{]}$
$=\frac{1}{2}\nabla\cdot\Big{[}\nabla\mathcal{G}(\tilde{\mathbf{x}})\,q^{*}(\tilde{\mathbf{x}})\Big{]}-\frac{1}{2}\nabla\cdot\Big{[}\nabla\mathcal{G}(\tilde{\mathbf{x}})\,q^{*}(\tilde{\mathbf{x}})\Big{]}$
@@
Consequently $q^{*}(\tilde{\mathbf{x}})$ remains in the kernel of the ~$\tilde{\mathcal{L}}^{*}_{\text{corr}}$, showing that this corrected time-rescaled SDE preserves the original invariant distribution.
@@
To verify the effectiveness of our time-rescaling approach, we analyzed the average number of iterations required to reach the desired success criterion during reward-guided Langevin dynamics on ORIBENCH-Single dataset. As shown in Tab. 3, Origen requires an average of 14.2 iterations to reach the desired reward level, whereas Origen reduced this to 12.8 iterations, demonstrating improved convergence speed without image quality degradation. Notably, the computational overhead introduced by adaptive rescaling is negligible, as the term $\nabla\log\mathcal{G}(\hat{\mathcal{R}}_{i})$ in Alg. 1 in the main paper can be efficiently computed via the chain rule, using the precomputed reward gradient ^$\nabla\hat{\mathcal{R}}_{i}$.
@@ Appendix F Implementation Details
For the Zero-1-to-3 [17] baseline, we generated the base image using FLUX-Schnell (512$\times$512) with 4 steps and encouraged the model to generate front-facing objects by adding the “facing front" prompt. The model was then conditioned on the target azimuth while keeping the polar angle fixed at 0° to generate novel foreground views, which were later composited with the background. The foreground was segmented using SAM [76], and missing background pixels were inpainted using LaMa [77] before composition.
@@
We empirically observed that our base model struggled to generate back-facing images when the prompt did not explicitly contain view information. We hypothesized that this issue arises because high-reward samples lie within an extremely sparse region of the conditional probability space, making them inherently challenging to sample, even when employing our proposed approach. To address this issue, we explicitly added the phrase “facing back" in facing back cases (i.e., where 90 $<$ $\phi_{i}^{\textit{az}}$ $<$ 270) and optimized the noise accordingly. With this simple adjustment, the model effectively generated both facing front and back images while maintaining high-quality outputs.
@@ Appendix I Computational Costs
In Tab. 7, we compare the computational costs of all baselines and Origen in terms of inference time and GPU memory consumption. Our base model (FLUX-Schnell [21]) and all guided generation methods were evaluated for generating 512$\times$512 images, while other methods followed their default settings. Guided generation methods are generally slower due to the need for reward backpropagation. All methods were evaluated under a setup of 50 Number of Function Evaluations (NFEs). Further analysis demonstrating the fast convergence of our reward-guided Langevin Dynamics is provided in Appendix C.2, suggesting that the computational cost can be partially reduced by using fewer NFEs. All measurements were conducted on a single A100 GPU with 80GB of memory.
@@
Appendix J More Qualitative Comparisons on Single Object Orientation Grounding
@@
Appendix K More Qualitative Comparisons on Multi Object Orientation Grounding

### v3

@@ 2 Related Work
For controlling image generation, ControlNet [28] and IP-Adapter [47] are commonly used to utilize a pretrained generative model to control for various conditions, though they require training data. For 3D orientation grounding, no public real-world training data is available, and collecting such data would be particularly challenging, especially for multi-object grounding, due to the need for diversity. To address this, recent works have introduced RL-based reward fine-tuning approaches [48, 49, blacktraining, 50, 51]. However, these methods require substantial computational resources and, more importantly, cause significant deviation from the pretrained data distribution when trained with task-specific rewards, leading to degraded image quality and reduced diversity [52]. Hence, instead of fine-tuning, we propose a test-time reward-guided framework that leverages a discriminative foundational model for guidance.
@@ 3.1 Problem Definition and Overview
Our goal is to achieve 3D orientation grounding in text-to-image generation using a one-step generative flow model [53] $\mathcal{F}_{\theta}$, based on test-time guidance without fine-tuning the model. Let $\mathbf{I}=\mathcal{F}_{\theta}(\mathbf{x},c)$ denote an image generated by $\mathcal{F}_{\theta}$ given an input text $c$ and a latent $\mathbf{x}$ sampled from a prior distribution $q=\mathcal{N}(\mathbf{0},\mathbf{I})$. The input text $c$ prompts the generation of an image containing a set of desired objects (e.g., “A person in a brown suit is directing a dog”). For each of the $N$ objects that appear in $c$, we associate a set of object phrases $\mathcal{W}=\{w_{i}\}_{i=1}^{N}$ (e.g., $\{\textrm{“\emph{dog}", “\emph{person}"}\}$) and a corresponding set of 3D orientation grounding conditions $\Phi=\{\phi_{i}\}_{i=1}^{N}$. Following the convention [20], each object orientation $\phi_{i}$ is represented by three Euler angles $\phi_{i}=\{\phi_{i,j}\}_{j=1}^{3}$: azimuth $\phi_{i,1}\in[0,360)$, polar angle $\phi_{i,2}\in[0,180)$, and rotation angle $\phi_{i,3}\in[0,360)$. The goal of 3D orientation grounding is to generate an image $\mathbf{I}$ in which the $N$ objects appear with their corresponding orientations $\Phi=\{\phi_{i}\}_{i=1}^{N}$.
@@
In particular, given a reward function $\mathcal{R}$ defined based on GroundingDINO and OrientAnything (which we will discuss in detail in Sec. 3.2), our goal is to find a latent sample $\mathbf{x}$ that maximizes the reward of its corresponding image, expressed as $\mathcal{R}(\mathcal{F}_{\theta}(\mathbf{x},c))$. For simplicity, we define the pullback of the reward function as $\hat{\mathcal{R}}=\mathcal{R}\circ\mathcal{F}_{\theta}$ and use this notation throughout.
@@ 3.3 Reward-Guided Langevin Dynamics
Reward-Guided Langevin Dynamics. Let $q=\mathcal{N}(\mathbf{0},\mathbf{I})$ denote the prior distribution, $\hat{\mathcal{R}}(\mathbf{x})$ be the pullback of a differentiable reward function, and $\mathbf{w}_{t}$ denote the standard Wiener process. As $t\rightarrow\infty$, the stationary distribution of the following Langevin dynamics
@@ 3.4 Reward-Adaptive Time Rescaling
[…]
3:for $i=0$ to M1M-1 do
4:  $\hat{\mathcal{R}}_{i}=\mathcal{R}(\mathcal{F}_{\theta}(\mathbf{x}_{i},c),\mathcal{W},\Phi)$ $\triangleright$ Reward Calc.
5:  $\gamma_{i}=\mathcal{G}(\hat{\mathcal{R}}_{i})\gamma$ $\triangleright$ Timestep rescaling
@@
7:  $\mathbf{x}_{i+1}=\sqrt{1-\gamma_{i}}\mathbf{x}_{i}+\gamma_{i}\eta\nabla\hat{\mathcal{R}}_{i}$$\frac{1}{2}\gamma_{i}\nabla\log\mathcal{G}(\hat{\mathcal{R}}_{i})$
8:       + $\sqrt{\gamma_{i}}\epsilon_{i}$ $\triangleright$ Update step
9:  if $\hat{\mathcal{R}}_{i}>\mathcal{R}_{*}$ then
10:   $\mathcal{R}_{*}=\hat{\mathcal{R}}_{i},\mathbf{x}_{*}=\mathbf{x}_{i}$
@@ 4 Experiments
Orientation Zero-1-to-3 C3DW FreeDoM ReNO Origen (Ours)
@@ 4.2 Evaluation Metrics
[…]
@@ 4.7 Computation Time
In our experiments, we set the number of iterations in our method to match the denoising steps in multi-step generative models (e.g., 50), resulting in comparable computation time—approximately 52.7 seconds per image. A detailed comparison is provided in Appendix I. Notably, for cases achieving angular accuracy within a tolerance of $\pm 22.5^{\circ}$ (Tab. 3), the average number of iterations to reach this threshold was 12.8, corresponding to an average of 13.5 seconds. Our approach can also serve as a post-refinement method for existing models, potentially further reducing computation time.
@@ 5 Conclusion
Acknowledgement
This work was supported by the NRF of Korea (RS-2023-00209723); IITP grants (RS-2022-II220594, RS-2023-00227592, RS-2024-00399817, RS-2025-25441313, RS-2025-25443318, RS-2025-02653113); the National Program for Excellence in SW, supervised by the IITP; and the Technology Innovation Program (RS-2025-02317326), all funded by the Korean government (MSIT and MOTIE), as well as by the DRB-KAIST SketchTheFuture Research Center.
@@ Appendix A Analysis on Norm-Based Regularization
where $\hat{\mathcal{R}}=\mathcal{R}\circ\mathcal{F}_{\theta}$ is the pullback of the reward function $\mathcal{R}$ and $y(\|\cdot\|)$ represents the probability density of a $\chi^{d}$ distribution.
[…]
@@ C.1 Non-Convergence of Direct Time-Rescaled SDE
[…]
@@ C.2 Time-Rescaled SDE with Correction Drift Term
The corresponding Fokker-Planck operator $\tilde{\mathcal{L}}^{*}_{\text{corr}}$ for this corrected time-rescaled SDE now annihilates the original stationary distribution $q^{*}(\mathbf{x})$:
[…]
@@
[…]
@@
To verify the effectiveness of our time-rescaling approach, we analyzed the average number of iterations required to reach the desired success criterion during reward-guided Langevin dynamics on ORIBENCH-Single dataset. As shown in Tab. 3, Origen requires an average of 14.2 iterations to reach the desired reward level, whereas Origen reduced this to 12.8 iterations, demonstrating improved convergence speed without image quality degradation. Notably, the computational overhead introduced by adaptive rescaling is negligible, as the term $\nabla\log\mathcal{G}(\hat{\mathcal{R}}_{i})$ in Alg. 1 in the main paper can be efficiently computed via the chain rule, using the precomputed reward gradient $\nabla\hat{\mathcal{R}}_{i}$.
@@ Appendix F Implementation Details
For the Zero-1-to-3 [17] baseline, we generated the base image using FLUX-Schnell ($\times512$) with 4 steps and encouraged the model to generate front-facing objects by adding the “facing front" prompt. The model was then conditioned on the target azimuth while keeping the polar angle fixed at 0° to generate novel foreground views, which were later composited with the background. The foreground was segmented using SAM [75], and missing background pixels were inpainted using LaMa [76] before composition.
@@
We empirically observed that our base model struggled to generate back-facing images when the prompt did not explicitly contain view information. We hypothesized that this issue arises because high-reward samples lie within an extremely sparse region of the conditional probability space, making them inherently challenging to sample, even when employing our proposed approach. To address this issue, we explicitly added the phrase “facing back" in facing back cases (i.e., where 90 << $\phi_{i}^{\textit{az}}$ << 270) and optimized the noise accordingly. With this simple adjustment, the model effectively generated both facing front and back images while maintaining high-quality outputs.
@@ Appendix I Computational Costs
In Tab. 7, we compare the computational costs of all baselines and Origen in terms of inference time and GPU memory consumption. Our base model (FLUX-Schnell [21]) and all guided generation methods were evaluated for generating $\times512$ images, while other methods followed their default settings. Guided generation methods are generally slower due to the need for reward backpropagation. All methods were evaluated under a setup of 50 Number of Function Evaluations (NFEs). Further analysis demonstrating the fast convergence of our reward-guided Langevin Dynamics is provided in Appendix C.2, suggesting that the computational cost can be partially reduced by using fewer NFEs. All measurements were conducted on a single A100 GPU with 80GB of memory.
Appendix J Extended Analysis of the Reward-Adaptive Monitor Function
Table 8: Extended analysis on the reward-adaptive monitor function. We ablate (i) the slope parameter $k$ and (ii) the step-size bounds $(s_{\min},s_{\max})$ to examine their effects on the sampling performance. For (i), $(s_{\min},s_{\max})$ are fixed to their default values of $(\tfrac{1}{3},\tfrac{4}{3})$, and for (ii), $k$ is fixed to its default value of $0.3$. Results indicate that overall performance is robust to variations in these hyperparameters. Best and second-best values are highlighted in bold and underlined, respectively.
(i) Ablation on slope parameter $k$ (ii) Ablation on step-size bounds ($s_{\min},s_{\max}$)
$k$ Azi. Acc. $\uparrow$ Azi. Err. $\downarrow$ CLIP $\uparrow$ VQA $\uparrow$ PickScore $\uparrow$     $s_{\min}$ $s_{\max}$ Azi. Acc. $\uparrow$ Azi. Err. $\downarrow$ CLIP $\uparrow$ VQA $\uparrow$ PickScore $\uparrow$
0.1 0.861 17.93 0.267 0.732 0.226     0.2 1.6 0.881 16.72 0.264 0.729 0.224
0.2 0.863 17.56 0.264 0.729 0.224     0.2 1.2 0.871 17.64 0.264 0.732 0.225
0.3 0.871 17.41 0.265 0.735 0.224     0.333 1.333 0.871 17.41 0.265 0.735 0.224
0.4 0.882 17.17 0.264 0.731 0.225     0.4 1.2 0.887 16.94 0.266 0.734 0.225
0.5 0.876 17.54 0.264 0.732 0.224     0.6 1.6 0.884 17.14 0.265 0.733 0.224
To further understand the effect of our reward-adaptive monitor function introduced in Sec. 3.4, we provide an extended analysis of its hyperparameters and their influence on model performance. As defined in Eq. 9, the monitor function adaptively modulates the sampling step size based on the reward, with three key hyperparameters: (i) $k$ controlling the slope (how fast the step size decreases as the reward increases) and (ii) $(s_{\min},s_{\max})$ defining the lower and upper bounds of the adaptive step size range.
We systematically vary these hyperparameters to evaluate the sensitivity of our sampling dynamics and its robustness to different configurations. Unless otherwise specified, all other hyperparameters are kept at their default values as reported in the main paper.
As shown in Tab. 8, our performance remains robust across all hyperparameter settings, with only negligible variations in the evaluation metrics. The default values were selected through a coarse grid search rather than fine-tuning, indicating that Origen does not rely on fragile hyperparameter choices. Moreover, all variants with the reward-adaptive monitor further outperform the fixed-step baseline reported in Tab. 1 of the main paper (denoted as Origen*), demonstrating its overall effectiveness.
Appendix K More Comprehensive Comparisons with Training-Based Methods
To complement the training-based results presented in Tab. 1 of the main paper, we provide more comprehensive quantitative comparisons with existing training-based approaches on the ORIBENCH-Single benchmark. Detailed descriptions of the baselines are provided below, and the quantitative results are summarized in the subsequent subsection.
K.1 Baselines.
Zero-1-to-3. [17] Please refer to ‘About Zero-1-to-3’ section in Appendix F.
C3DW. [18] Please refer to ‘About C3DW’ section in Appendix F.
DDPO. [77] We evaluated RL-based fine-tuning method for orientation alignment. Specifically, we adopt DDPO [77], which fine-tunes a diffusion model via policy gradient updates on a learned reward signal. We note that existing RL-based diffusion fine-tuning methods are designed for either unconditional or text-conditional generation  [48, 49, 77]. Thus, to enable orientation conditioning while leveraging existing code of DDPO, we opt to appending orientation phrases to the input text prompts (e.g., “object is viewed from front-right, low-angle (azimuth 315°, polar 61°, rotation 90°)”). To fine-tune with our orientation grounding reward, we constructed a single-object text–orientation paired training dataset using 50k orientations from the OrientAnything [20] training dataset and captions from the Cap3D dataset  [78, 79]. For other fine-tuning setups (e.g., hyperparameters), we primarily followed the default configuration of original DDPO to ensure fair comparisons.
SV3D. [80] SV3D is a multi-view diffusion model generates images conditioned on input image and target view angle. SV3D generates 21 views simultaneously, and we used the one generated image with target azimuth. The whole pipeline used in evaluation is same as Zero-1-to-3 [17], so please refer to ‘About Zero-1-to-3’ section in Appendix F.
MV-Adapter. [81] MV-Adapter is a multi-view diffusion model generates images conditioned on input image and target view angle. MV-Adapter generates 6 views simultaneously, and we used the one generated image with target azimuth. The whole pipeline used in evaluation is same as Zero-1-to-3 [17], so please refer to ‘About Zero-1-to-3’ section in Appendix F.
K.2 Results.
As shown in Table 9, Origen achieves substantially higher orientation alignment accuracy than all training-based baselines while maintaining competitive text-to-image alignment performance. These results demonstrate that our inference-time reward-guided sampling consistently outperforms training-based methods that rely on explicit retraining or fine-tuning for orientation control.
Table 9: Quantitative comparisons on Training-based Methods. Origen maintains high accuracy even when compared with more recent training-based methods. Best and second-best results are highlighted in bold and underlined, respectively.
Id Model Orientation Alignment Text-to-Image Alignment
Azi, Acc. @22.5° $\uparrow$ Azi, Abs. Err. $\downarrow$ CLIP $\uparrow$ VQA $\uparrow$ PickScore $\uparrow$
1 Zero-1-to-3 [17] 0.499 59.03 0.272 0.663 0.213
2 C3DW [18] 0.426 64.77 0.220 0.439 0.197
3 DDPO [77] 0.494 40.83 0.256 0.702 0.233
4 SV3D [80] 0.410 60.26 0.274 0.313 0.196
5 MV-Adapter [81] 0.486 50.19 0.204 0.313 0.196
6 Origen (Ours) 0.871 17.41 0.265 0.735 0.224
Appendix L Reward-Guided Langevin Dynamics for Other Rewards
To assess the broader applicability of our reward-guided Langevin dynamics sampling, which is inherently reward-agnostic, we extended our experiments beyond 3D orientation grounding to additional tasks, including layout-to-image and depth-guided generation.
Table 10: Quantitative comparisons on layout-to-image and depth-guided generation. Best and second-best results are highlighted in bold and underlined, respectively.
Layout-to-Image Generation Depth-Guided Generation
Id Model Spatial Alignment Text-to-Image Alignment Depth Alignment Text-to-Image Alignment
  mIoU $\uparrow$  HRS $\uparrow$  CLIP $\uparrow$ VQA $\uparrow$  PickScore $\uparrow$  AbsRel $\downarrow$  CLIP $\uparrow$  VQA $\uparrow$  PickScore $\uparrow$
1 FreeDoM [35] 0.244 60 0.261 0.640 0.261 6.34 0.256 0.672 0.223
2 ReNO [22] 0.287 70 0.261 0.655 0.229 6.67 0.265 0.705 0.233
3 ORIGEN (Ours) 0.344 72 0.284 0.659 0.229 4.62 0.264 0.686 0.230
L.1 Experimental Setup
We used FreeDoM [35] and ReNO [22] as baselines to highlight the effectiveness of our sampling strategy.
Layout-to-Image Generation. For evaluation setup, we randomly sampled 100 examples from the HRS-Spatial dataset [82] and used GroundingDINO [54] as the reward model to guide object positioning based on bounding-box conditions. We evaluated the results using mIoU (measured by other detection model [83]) and HRS score.
Depth-Guided Generation. For evaluation setup, we randomly sampled 100 image-caption pairs from the ORIBENCH-Single dataset and generated pseudo-ground-truth depth maps using DepthAnything-V2 [84]. We then used DepthAnything-V2 as a reward model to guide image generation and evaluated the results using Absolute Relative Error (AbsRel).
Condition FreeDoM [35] ReNO [22] Origen (Ours)
Layout-to-Image Refer to caption
Depth-Guided Refer to caption
Figure 9: Qualitative comparisons on layout-to-image (rows 1-2) and depth-guided generation (rows 3-4). Origen achieves better condition alignment and overall image quality compared to other guided-generation methods [35, 22].
L.2 Results
As shown in Fig. 9 and Tab. 10, Origen consistently outperforms other guided-generation methods [35, 22] in condition alignment and image quality across both layout-to-image and depth-guided generation tasks. These results demonstrate that our method generalizes effectively to diverse reward alignment tasks.
@@
Appendix M More Qualitative Comparisons on Single Object Orientation Grounding
@@
Appendix N More Qualitative Comparisons on Multi Object Orientation Grounding